OpenStudy (dangerousjesse):

How many pounds are in 0.0379 g?

OpenStudy (anonymous):

Dimensional analysis/conversion factors

OpenStudy (dangerousjesse):

I know for a fact that it's \(\large 8.34\times e^{-5} lb\), but I just want to double check.

OpenStudy (anonymous):

google says \(\huge \frac{1gram}{0.00220462lb}\), following: \(\huge \frac{0.00220462lb}{\cancel{1g}}*\frac{0.0379 \cancel{g}}{1}\)

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} grams&lbs \\\hline\\ 453.592&1\\ 0.0379&x \end{array}\implies \cfrac{453.592}{ 0.0379 }=\cfrac{1}{x}\)

OpenStudy (anonymous):

8.35552e-5 = 8.35552 * \(\large 10^{-5}\) = 0.0000835552